# FDS: Feedback-guided Domain Synthesis with Multi-Source Conditional Diffusion Models for Domain Generalization

### v2

@@ Abstract
Domain Generalization techniques aim to enhance model robustness by simulating novel data distributions during training, typically through various augmentation or stylization strategies. However, these methods frequently suffer from limited control over the diversity of generated images and lack assurance that these images span distinct distributions. To address these challenges, we propose FDS, Feedback-guided Domain Synthesis, a novel strategy that employs diffusion models to synthesize novel, pseudo-domains by training a single model on all source domains and performing domain mixing based on learned features. By incorporating images that pose classification challenges to models trained on original samples, alongside the original dataset, we ensure the generation of a training set that spans a broad distribution spectrum. Our comprehensive evaluations demonstrate that this methodology sets new benchmarks in domain generalization performance across a range of challenging datasets, effectively managing diverse types of domain shifts. The implementation is available at: https://github.com/Mehrdad-Noori/FDS.git.
@@ 1 Introduction
Deep learning architectures, including Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), have significantly advanced the field of computer vision, achieving state-of-art results in tasks like classification, semantic segmentation, and object detection. Despite these advancements, such models commonly operate under the simplistic assumption that training data (source domain) and post-deployment data (target domain) share identical distributions. This overlook of distributional shifts results in performance degradation when models are exposed to out-of-distribution (OOD) data [1, 2]. Domain adaptation (DA) [3, 4, 5] and Test-Time Adaptation (or Training) [6, 7, 8] strategies have been developed to mitigate this issue by adjusting models trained on the source domain to accommodate a predefined target domain. Nonetheless, these strategies are constrained by their dependence on accessible target domain data for adaptation, a prerequisite that is not always feasible in real-world applications. Furthermore, adapting models to each novel target domain entails considerable computational overhead, presenting a practical challenge to their widespread implementation.
Domain generalization (DG) [9] aims to solve the issue of domain shift by training models using data from one or more source domains, so they perform well out-of-the-box on new, unseen domains. Recently, various techniques have been developed to tackle this problem [10, 11], including domain aligning [12, 13, 14], meta-learning [15, 16], data augmentation [17, 18, 19], ensemble learning[20], self-supervised learning [21, 22], and regularization methods [23, 24]. These strategies are designed to make models more adaptable and capable of handling data that they were not explicitly trained on, making them more useful in real-world situations where the exact nature of future data cannot be predicted. Among these techniques, a notable category focuses on synthesizing samples from different distributions to mimic target distributions. This is achieved through strategies like image transformation [17, 25], style transfer [26, 27], learnable augmentation networks [28, 29, 30], and feature-level stylization [20, 31]. However, many of these methods face challenges in controlling the synthesis process, often resulting in limited diversity where primarily only textures are altered.
@@
In this study, we introduce an innovative approach using diffusion model, named Feedback-guided Domain Synthesis (FDS), to address the challenge of domain generalization. Known for their exceptional ability to grasp intricate distributions and semantics, diffusion models excel at producing high-quality, realistic samples [32, 33, 34, 35]. We exploit this strength by training a single diffusion model that is conditioned on various domains and classes present in the training dataset, aiming to master the distribution of source domains. As illustrated in Figure 1, through the process of domain interpolation and mixing during generation, we create images that appear to originate from novel, pseudo-domains. To ensure the development of a robust classifier, we initially select generated samples that are difficult for a model trained solely on the original source domains to classify. Subsequently, we train the model using a combination of these challenging images and the original dataset. In this manner, we ensure that our model is fed with images of the widest possible diversity, thereby significantly enhancing its ability to generalize across unseen domains. Our contributions can be summarized as follows:
@@ 2.2 Diffusion Models
Despite the capabilities of diffusion models in generating images, their application in domain generalization has been minimally explored. Yue et al. [54] proposed a method called DSI to address OOD prediction by transforming testing samples back to the training distribution using multiple diffusion models, each trained on a single source distribution. While effective, DSI requires multiple models during prediction, making it impractical for real-time deployment. In contrast, our method only uses diffusion during training to synthesize pseudo-domains and select more challenging images, keeping the computational complexity the same during testing while achieving significantly better performance.
In another study, CDGA [55] uses diffusion models to generate synthetic images that fill the gap between different domain pairs through a simple interpolation. While effective, CDGA relies on generating a large number of synthetic images (e.g., 5M for PACS) and employs naive interpolation. Additionally, CDGA generates samples with additional text descriptions that are not provided in standard DG benchmarks. In contrast, our method employs novel and more efficient interpolation techniques and a unique filtering mechanism that selects challenging images. We demonstrate that both of these innovations are crucial and significantly effective for improving generalization.
@@ 3 Theoretical Motivation
Here, $|\mathcal{D}^{i}|$ counts the number of samples in domain $i$, with $x_{i}^{k}$ as the $k$-th sample and $y_{i}^{k}$ its label. However, the accuracy of ERM-trained models drops when data shifts occur across domains due to inadequate OOD generalization. To enhance ERM, Chapelle et al. introduced Vicinal Risk Minimization (VRM) [57], which substitutes point-wise estimates with density estimation in the vicinity distribution around each observation within each domain. Practical implementation often involves data augmentation to introduce synthetic samples from these density estimates. Traditional augmentation processes a single data point $\tilde{x}_{i}^{k}$ from domain $i$ to yield $\tilde{x}_{i}^{k}=g(\tilde{x}_{i}^{k})$, where $g(\cdot)$ denotes a basic transformation and $\tilde{x}_{i}^{k}$ is the modified data point.
@@ 4 Method
To advance the generalization ability of our classifier, we aim to utilize generated image samples that traverse domain gaps yet retain class semantics. This objective is realized through a three-step methodology of our proposed method FDS. Step 1: we begin by training an image generator that masters the class-specific distributions of all source domains, enabling the synthesis of class-consistent samples within those domains. Step 2: we then introduce a strategy for generating inter-domain images to span the domain gaps effectively. Step 3: last, we filter overly simplistic samples from synthetic inter-domain images and train the final model on this refined set alongside original domain images for enhanced OOD generalization. The process is illustrated in Figure 2.
@@ 4.2.2 Condition Level Interpolation
The core of this strategy involves blending these condition representations using a mix coefficient $\alpha$, leading to a unified condition:
@@ 4.3 Filtering Mechanism
Let $\mathcal{C}_{i,j}^{k}$ be the subset of samples in $S_{i,j}^{k}$ with are correctly classified by $h(x)$:
@@ 5 Experimental Setup
Table 1: Leave-one-out accuracy (%) results on the PACS, VLCS, and OfficeHome benchmarks. "Aug." indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.
@@
Method Aug. PACS VLCS Office  Avg. Standard Methods ERM (baseline[44] 85.5 $\pm$0.2 77.5 $\pm$0.4 66.5 $\pm$0.3 76.5 ERM (reproduced) 84.3 $\pm$1.1 76.2 $\pm$1.1 64.6 $\pm$1.1 75.0 IRM [60] 83.5 $\pm$0.8 78.5 $\pm$0.5 64.3 $\pm$2.2 75.4 GroupDRO [47] 84.4 $\pm$0.8 76.7 $\pm$0.6 66.0 $\pm$0.7 75.7 Mixup [61] 84.6 $\pm$0.6 77.4 $\pm$0.6 68.1 $\pm$0.3 76.7 CORAL [62] 86.2 $\pm$0.3 78.8 $\pm$0.6 68.7 $\pm$0.3 77.9 MMD [63] 84.6 $\pm$0.5 77.5 $\pm$0.9 66.3 $\pm$0.1 76.1 DANN [64] 83.6 $\pm$0.4 78.6 $\pm$0.4 65.9 $\pm$0.6 76.0 SagNet [65] 86.3 $\pm$0.2 77.8 $\pm$0.5 68.1 $\pm$0.1 77.4 RSC [23] 85.2 $\pm$0.9 77.1 $\pm$0.5 65.5 $\pm$0.9 75.9 Mixstyle [20] 85.2 $\pm$0.3 77.9 $\pm$0.5 60.4 $\pm$0.3 74.5 mDSDI [66] 86.2 $\pm$0.2 79.0 $\pm$0.3 69.2 $\pm$0.4 78.1 SelfReg [41] 85.6 $\pm$0.4 77.8 $\pm$0.9 67.9 $\pm$0.7 77.1 DCAug [59] 86.1 $\pm$0.7 78.6 $\pm$0.4 68.3 $\pm$0.4 77.7 CDGA [55] 88.5 $\pm$0.5 79.6 $\pm$0.3 68.2 $\pm$0.6 78.8 ERM + FDS (ours) 88.8 $\pm$0.1 79.8 $\pm$0.5 71.1 $\pm$0.1 79.9 WA Methods SWAD (baseline[24] 88.1 $\pm$0.1 79.1 $\pm$0.1 70.6 $\pm$0.2 79.3 SWAD (reproduced) 88.1 $\pm$0.4 78.9 $\pm$0.5 70.3 $\pm$0.4 79.1 SelfReg SWA [41] 86.5 $\pm$0.3 77.5 $\pm$0.0 69.4 $\pm$0.2 77.8 DNA [67] 88.4 $\pm$0.1 79.0 $\pm$0.1 71.2 $\pm$0.1 79.5 DIWA [68] 88.8 $\pm$0.4 79.1 $\pm$0.2 71.0 $\pm$0.1 79.6 TeachDCAug [59] 88.4 $\pm$0.2 78.8 $\pm$0.4 70.4 $\pm$0.2 79.2 SWAD + FDS (ours) 90.5 $\pm$0.3 79.7 $\pm$0.5 73.5 $\pm$0.4 81.3
@@
Datasets.  Following [44], we compare our proposed approach to the current state-of-art using three challenging datasets - 12 individual target domains - with different characteristics: PACS [42], VLCS [69], and OfficeHome [70]. The PACS dataset has a total of 9,991 photos divided into four distinct domains, $d\in$Art, Cartoon, Photo, Sketch} and seven distinct classes. The second dataset, VLCS, comprises 10,729 photos from four separate domains, $d\in$Caltech101, LabelMe , SUN09, VOC2007} and five different classes. The third dataset, OfficeHome, includes a total of 15,588 photos taken from four domains, $d\in$Art, Clipart, Product, Real}, and 65 classes.
@@
Implementation Details.  To ensure a fair comparison, we adopt the DomainBed framework [44], a comprehensive benchmark that encompasses prominent domain generalization (DG) methodologies under a uniform evaluation protocol. Following this framework, we employ a leave-one-out strategy for DG dataset assessment where one domain serves as the test set while the others form the training set. A subset of the training data, constituting 20%, is designated as the validation set [^1] The model with peak accuracy on this validation set is selected for evaluation on the test domain, providing unseen domain accuracy.. The aggregate result for each dataset represents the mean accuracy derived from varying the test domain. To ensure reliability, experiments are replicated three times, each with a unique seed. Moreover, to rigorously test our method, we try both ERM and SWAD classifiers with FDS. The former is considered a baseline in standard training methods, while the latter serves as a baseline for weight averaging (WA) methods. SWAD is essentially ERM but with weight averaging applied during training using multiple steps based on the validation set. To have a fair comparison with other methods, for the ERM baseline, we strictly follow the hyperparameter tuning proposed in [24]. For SWAD, as in the original work, we did not tune any parameters and used the default values. For image synthesis, the original Stable Diffusion framework [34] is utilized with $\mathrm{DDIM}\!=\!50$ steps. The PACS and VLCS datasets prompt the generation of $N\!=\!32,000$ samples per class, whereas OfficeHome, with its 65 classes, necessitates $N\!=\!16,000$ samples. Image generation spans an interpolation range of $\alpha\!\in\![0.3,0.7]$ and a Noise Level Interpolation range of $\mathcal{T}\!\in\![20,45]$. This diversified parameter selection, rather than optimizing hyperparameters per dataset, acknowledges each domain’s unique shift. Our filtering mechanism then identifies the most informative images from the generated pool. The selection of an optimal dataset size, $N_{L}$, underwent extensive exploration, detailed further in Section 6.2.
@@ 6.1 Comparison with the State-of-the-art
Furthermore, when comparing with previous SOTA methods, our method outperforms CDGA [55], by 0.3% on PACS, 0.2% on VLCS, and 2.9% on OfficeHome. In the context of weight averaging methods, our approach surpasses DIWA [68], which trains multiple independent models, by 1.7% on PACS, 0.6% on VLCS, and 2.5% on OfficeHome, setting a new benchmark for domain generalization. Please refer to the Supplementary Material for the full, detailed results for each dataset and its domains.

### v3

@@ Abstract
Domain Generalization techniques aim to enhance model robustness by simulating novel data distributions during training, typically through various augmentation or stylization strategies. However, these methods frequently suffer from limited control over the diversity of generated images and lack assurance that these images span distinct distributions. To address these challenges, we propose FDS, Feedback-guided Domain Synthesis, a novel strategy that employs diffusion models to synthesize novel, pseudo-domains by training a single model on all source domains and performing domain mixing based on learned features. By incorporating images that pose classification challenges to models trained on original samples, alongside the original dataset, we ensure the generation of a training set that spans a broad distribution spectrum. Our comprehensive evaluations demonstrate that this methodology sets new benchmarks in domain generalization performance across a range of challenging datasets, effectively managing diverse types of domain shifts. The code can be found at: https://github.com/Mehrdad-Noori/FDS.
@@ 1 Introduction
Deep learning architectures, including Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), have significantly advanced the field of computer vision, achieving state-of-art results in tasks like classification, semantic segmentation, and object detection. Despite these advancements, such models commonly operate under the simplistic assumption that training data (source domain) and post-deployment data (target domain) share identical distributions. This overlook of distributional shifts results in performance degradation when models are exposed to out-of-distribution (OOD) data [1, 2]. Domain adaptation (DA) [3, 4] and Test-Time Adaptation (or Training) [5, 6, 7] strategies have been developed to mitigate this issue by adjusting models trained on the source domain to accommodate a predefined target domain. Nonetheless, these strategies are constrained by their dependence on accessible target domain data for adaptation, a prerequisite that is not always feasible in real-world applications. Furthermore, adapting models to each novel target domain entails considerable computational overhead, presenting a practical challenge to their widespread implementation.
Domain generalization (DG) [8] aims to solve the issue of domain shift by training models using data from one or more source domains so that they perform well out-of-the-box on new, unseen domains. Recently, various techniques have been developed to tackle this problem [9, 10], including domain aligning [11, 12, 13], meta-learning [14, 15], data augmentation [16, 17, 18], ensemble learning[19], self-supervised learning [20, 21], and regularization methods [22, 23]. These strategies are designed to make models more adaptable and capable of handling data that they were not explicitly trained on, making them more useful in real-world situations where the exact nature of future data cannot be predicted. Among these techniques, a notable category focuses on synthesizing samples from different distributions to mimic target distributions. This is achieved through strategies like image transformation [16, 24], style transfer [25, 26], learnable augmentation networks [27, 28, 29], and feature-level stylization [19, 30, 31]. However, many of these methods face challenges in controlling the synthesis process, often resulting in limited diversity where primarily only textures are altered.
@@
In this study, we introduce an innovative approach using diffusion model, named Feedback-guided Domain Synthesis (FDS), to address the challenge of domain generalization. Known for their exceptional ability to grasp intricate distributions and semantics, diffusion models excel at producing high-quality, realistic samples [32, 33, 34, 35]. We exploit this strength by training a single diffusion model that is conditioned on various domains and classes present in the training dataset, aiming to master the distribution of source domains.
As illustrated in Figure 1, through the process of domain interpolation and mixing during generation, we create images that appear to originate from novel, pseudo-domains. To ensure the development of a robust classifier, we initially select generated samples that are difficult for a model trained solely on the original source domains to classify. Subsequently, we train the model using a combination of these challenging images and the original dataset. In this manner, we make sure that our model is fed with images of the widest possible diversity, thereby significantly enhancing its ability to generalize across unseen domains. Our contributions can be summarized as follows:
@@ 2.2 Diffusion Models
Despite the capabilities of diffusion models in generating images, their application in domain generalization has been minimally explored. Miao et al. [54] introduced DomainDiff, a method that boosts OOD generalization by training a Word-to-Image Mapping (WIM) with diffusion models to generate additional synthetic data. Yue et al. [55] proposed a method called DSI to address OOD prediction by transforming testing samples back to the training distribution using multiple diffusion models, each trained on a single source distribution. While effective, DSI requires multiple models during prediction, making it impractical for real-time deployment. In contrast, our method only uses diffusion during training to synthesize pseudo-domains and select more challenging images, keeping the computational complexity the same during testing while achieving significantly better performance.
In another study, CDGA [56] uses diffusion models to generate synthetic images that fill the gap between different domain pairs through a simple interpolation. While effective, CDGA relies on generating a large number of synthetic images (e.g., 5M for PACS) and employs naive interpolation. Additionally, CDGA generates samples with additional text descriptions that are not provided in standard DG benchmarks. In contrast, our method employs novel and more efficient interpolation techniques and a unique filtering mechanism that selects challenging images. We demonstrate that both of these innovations are crucial for improving generalization.
@@ 3 Theoretical Motivation
Here, $|\mathcal{D}^{i}|$ counts the number of samples in domain $i$, with $x_{i}^{k}$ as the $k$-th sample and $y_{i}^{k}$ its label. However, the accuracy of ERM-trained models drops when data shifts occur across domains due to inadequate OOD generalization. To enhance ERM, Chapelle et al. introduced Vicinal Risk Minimization (VRM) [58], which substitutes point-wise estimates with density estimation in the vicinity distribution around every observation within each domain. Practical implementation often involves data augmentation to introduce synthetic samples from these density estimates. Traditional augmentation processes a single data point $x_{i}^{k}$ from domain $i$ to yield $\tilde{x}_{i}^{k}=g(x_{i}^{k})$, where $g(\cdot)$ denotes a basic transformation and $\tilde{x}_{i}^{k}$ is the modified data point.
@@ 4 Method
To advance the generalization ability of our classifier, we aim to utilize generated image samples that traverse domain gaps yet retain class semantics. This objective is realized through a three-step methodology at the core of our FDS approach. Step 1: we begin by training an image generator that masters the class-specific distributions of all source domains, enabling the synthesis of class-consistent samples within those domains. Step 2: we then introduce a strategy for generating inter-domain images to span the domain gaps effectively. Step 3: last, we filter overly simplistic samples from synthetic inter-domain images and train the final model on this refined set alongside original domain images for enhanced OOD generalization. The process is illustrated in Figure 2.
@@ 4.2.2 Condition Level Interpolation
The core of this strategy involves blending these condition representations using a mixing coefficient $\alpha$, leading to a unified condition:
@@ 4.3 Filtering Mechanism
Let $\mathcal{C}_{i,j}^{k}$ be the subset of samples in $S_{i,j}^{k}$ which are correctly classified by $h(x)$:
@@ 5 Experimental Setup
Method Aug. PACS VLCS Office  Avg. Standard Methods ERM (baseline[44] 85.5 $\pm$0.2 77.5 $\pm$0.4 66.5 $\pm$0.3 76.5 ERM (reproduced) 84.3 $\pm$1.1 76.2 $\pm$1.1 64.6 $\pm$1.1 75.0 IRM [61] 83.5 $\pm$0.8 78.5 $\pm$0.5 64.3 $\pm$2.2 75.4 GroupDRO [47] 84.4 $\pm$0.8 76.7 $\pm$0.6 66.0 $\pm$0.7 75.7 Mixup [62] 84.6 $\pm$0.6 77.4 $\pm$0.6 68.1 $\pm$0.3 76.7 CORAL [63] 86.2 $\pm$0.3 78.8 $\pm$0.6 68.7 $\pm$0.3 77.9 MMD [64] 84.6 $\pm$0.5 77.5 $\pm$0.9 66.3 $\pm$0.1 76.1 DANN [65] 83.6 $\pm$0.4 78.6 $\pm$0.4 65.9 $\pm$0.6 76.0 SagNet [66] 86.3 $\pm$0.2 77.8 $\pm$0.5 68.1 $\pm$0.1 77.4 RSC [22] 85.2 $\pm$0.9 77.1 $\pm$0.5 65.5 $\pm$0.9 75.9 Mixstyle [19] 85.2 $\pm$0.3 77.9 $\pm$0.5 60.4 $\pm$0.3 74.5 mDSDI [67] 86.2 $\pm$0.2 79.0 $\pm$0.3 69.2 $\pm$0.4 78.1 SelfReg [41] 85.6 $\pm$0.4 77.8 $\pm$0.9 67.9 $\pm$0.7 77.1 DCAug [60] 86.1 $\pm$0.7 78.6 $\pm$0.4 68.3 $\pm$0.4 77.7 DomainDiff [54] 85.6 $\pm$0.6 63.7 $\pm$0.6 DSI [55] 86.9 $\pm$1.4 CDGA [56] 88.5 $\pm$0.5 79.6 $\pm$0.3 68.2 $\pm$0.6 78.8 ERM + FDS (ours) 88.8 $\pm$0.1 79.8 $\pm$0.5 71.1 $\pm$0.1 79.9 WA Methods SWAD (baseline[23] 88.1 $\pm$0.1 79.1 $\pm$0.1 70.6 $\pm$0.2 79.3 SWAD (reproduced) 88.1 $\pm$0.4 78.9 $\pm$0.5 70.3 $\pm$0.4 79.1 SelfReg SWA [41] 86.5 $\pm$0.3 77.5 $\pm$0.0 69.4 $\pm$0.2 77.8 DNA [68] 88.4 $\pm$0.1 79.0 $\pm$0.1 71.2 $\pm$0.1 79.5 DIWA [69] 88.8 $\pm$0.4 79.1 $\pm$0.2 71.0 $\pm$0.1 79.6 TeachDCAug [60] 88.4 $\pm$0.2 78.8 $\pm$0.4 70.4 $\pm$0.2 79.2 SWAD + FDS (ours) 90.5 $\pm$0.3 79.7 $\pm$0.5 73.5 $\pm$0.4 81.3
Table 1: Leave-one-out accuracy (%) results on the PACS, VLCS, and OfficeHome benchmarks. “Aug.” indicates whether advanced augmentation or domain mixing techniques are used. The best results and second-best results are highlighted.
@@
Datasets.  Following [44], we compare our proposed approach to the current state-of-art using three challenging datasets–12 individual target domains–with different characteristics: PACS [42], VLCS [70], and OfficeHome [71]. The PACS dataset has a total of 9,991 photos divided into four distinct domains, $d\in$Art, Cartoon, Photo, Sketch}, and seven distinct classes. The second dataset, VLCS, comprises 10,729 photos from four separate domains, $d\in$Caltech101, LabelMe , SUN09, VOC2007}, and five different classes. The third dataset, OfficeHome, includes a total of 15,588 photos taken from four domains, $d\in$Art, Clipart, Product, Real}, and 65 classes.
@@
Implementation Details.  To ensure a fair comparison, we adopt the DomainBed framework [44], a comprehensive benchmark that encompasses prominent domain generalization (DG) methodologies under a uniform evaluation protocol. Following this framework, we employ a leave-one-out strategy for DG dataset assessment where one domain serves as the test set while the others form the training set. A subset of the training data, constituting 20%, is designated as the validation set111 The model with peak accuracy on this validation set is selected for evaluation on the test domain, providing unseen domain accuracy.. The aggregate result for each dataset represents the mean accuracy derived from varying the test domain. To ensure reliability, experiments are replicated three times, each with a unique seed. Moreover, to rigorously test our method, we try both ERM and SWAD classifiers with FDS. The former is considered a baseline in standard training methods, while the latter serves as a baseline for weight averaging (WA) methods. SWAD is essentially ERM but with weight averaging applied during training using multiple steps based on the validation set. To have a fair comparison with other methods, for the ERM baseline, we strictly follow the hyperparameter tuning proposed in [23]. For SWAD, as in the original work, we did not tune any parameters and used the default values. For image synthesis, the original Stable Diffusion framework [34] is utilized with $\mathrm{DDIM}\!=\!50$ steps. The PACS and VLCS datasets prompt the generation of $N\!=\!32,000$ samples per class, whereas OfficeHome, with its 65 classes, necessitates $N\!=\!16,000$ samples. Image generation spans an interpolation range of $\alpha\!\in\![0.3,0.7]$ and a Noise Level Interpolation range of $\mathcal{T}\!\in\![20,45]$. This diversified parameter selection, rather than optimizing hyperparameters per dataset, acknowledges each domain’s unique shift. Our filtering mechanism then identifies the most informative images from the generated pool. The impact of dataset size, $N_{L}$, is studied further in Section 6.2.
@@ 6.1 Comparison with the State-of-the-art
Furthermore, when comparing with previous SOTA methods, our method outperforms CDGA [56] by 0.3% on PACS, 0.2% on VLCS, and 2.9% on OfficeHome. In the context of weight averaging methods, our approach surpasses DIWA [69], which trains multiple independent models, by 1.7% on PACS, 0.6% on VLCS, and 2.5% on OfficeHome, setting a new benchmark for domain generalization. Please refer to the Supplementary Material for the full, detailed results for each dataset and its domains.
Module Target Domains Art Cartoon Photo Sketch Avg. Baseline (SWAD [23]) 89.49 $\pm$0.2 83.65 $\pm$0.4 97.25 $\pm$0.2 82.06 $\pm$1.0 88.11 $\pm$0.4 + Basic Gen. 89.87 $\pm$0.1 85.59 $\pm$0.6 97.50 $\pm$0.3 83.07 $\pm$0.4 89.01 $\pm$0.4 + Interpolation 91.38 $\pm$0.2 85.20 $\pm$0.6 97.73 $\pm$0.1 84.27 $\pm$0.9 89.65 $\pm$0.4 + Filtering 91.80 $\pm$0.3 86.03 $\pm$0.8 98.05 $\pm$0.2 86.11 $\pm$0.1 90.50 $\pm$0.3
Table 2: Comparative analysis of FDS component effects on accuracy (%) across PACS dataset domains. “Basic Gen.” refers to generation without interpolation or filtering.